A uniform meter stick (a ruler 100 cm long) is supported by a fulcrum at its 20 cm mark balances when a 200 g mass is suspended at the 0 cm end. Assume that the center of mass of the meter stick is exactly at its middle, the 50 cm mark. Determine the mass of the meter stick, in grams. Please show your work.
Answer #1

Suppose a meter stick. Now, its mass is concenterated on the 50 cm mark. Now, we know that it balances on 20 cm mark which is 30 cm to the left of central 50 cm mark.

Also, the 200 g mass is at 0 cm mark. which is 20 cm left to the fulcrum

Now, using definition of torque, we have

let x be the mass of the meter stick

200*20 = x * 30

x = 200*20 / 30

x = 133.33 gms


answered by: ANURANJAN SARSAM